Ibsen Company makes two products from a common input. Joint processing costs up to the split-off point total $51,000 a year. The company allocates these costs to the joint products on the basis of their total sales values at the split-off point. Each product may be sold at the split-off point or processed further. Data concerning these products appear below:

Product X Product Y Total
Allocated joint processing costs $ 20,400 $ 30,600 $ 51,000
Sales value at split-off point $ 20,000 $ 30,000 $ 50,000
Costs of further processing $ 23,300 $ 17,600 $ 40,900
Sales value after further processing $ 38,800 $ 56,500 $ 95,300

Required:

a. What is financial advantage (disadvantage) of processing Product X beyond the split-off point? (Negative amount should be indicated by a minus sign.)

b. What is financial advantage (disadvantage) of processing Product Y beyond the split-off point?

c. What is the minimum amount the company should accept for Product X if it is to be sold at the split-off point?

d. What is the minimum amount the company should accept for Product Y if it is to be sold at the split-off point?
Answer #1

a. Financial advantage (disadvantage) = Additional revenue - further processing cost

Product X = (38800-20000)-23300 = $-4500

b. Product Y = (56500-30000)-17690 = $

c. Minimum amount = Allocated joint processing cost = $20500

d. Minimum amount = $30600